# Dissipation-dispersion analysis of fully-discrete implicit discontinuous Galerkin methods and application to stiff hyperbolic problems

### v2

@@ Abstract
The application of discontinuous Galerkin (DG) schemes to hyperbolic systems of conservation laws requires a careful interplay between the space discretization, carried out with local polynomials and numerical fluxes at inter-cells, and a high-order time-integration to yield the final update. An important concern is how the scheme modifies the solution through the notions of numerical diffusion and dispersion. The analysis of these artifacts is well known for finite volume schemes, but it becomes more complex in the DG case. In particular, as far as we know, no analysis of this type has been considered for implicit integration with DG space discretization. The first part of this work intends to fill this gap, showing that the choice of the implicit Runge-Kutta scheme impacts deeply on the quality of the solution. We analyze dispersion-diffusion properties to select the best combination of the space-time discretization for high Courant numbers.
In the second part of this work, we apply our findings to the integration of stiff hyperbolic systems with DG schemes. Implicit time-integration schemes leverage superior stability properties enabling the selection of time-steps based solely on accuracy requirements, thereby bypassing the need for minute time-steps. High-order schemes require the introduction of local space limiters which make the whole implicit scheme highly nonlinear. To mitigate the numerical complexity of these schemes, we propose to use appropriate space limiters that can be precomputed on a first-order prediction of the solution. This approach follows the methodology proposed by Puppo et al. (Commun. Comput. Phys., 2024) for high-order finite volume schemes. Numerical experiments explore the performance of this technique on scalar equations and systems.
@@ 1 Introduction
From a numerical viewpoint, stiffness imposes severe restrictions on time-step sizes when using explicit time-integration methods, which significantly increases computational costs. In fact, the Courant-Friedrichs-Lewy (CFL) condition forces the time-step to be small enough to ensure stability, which in turn imposes accuracy on fast-moving waves. To better address the dynamics of the slower-moving fluid, implicit time-integration methods are often used. These methods are unconditionally stable and do not adhere to the CFL condition, allowing to choose the time-step based on the fluid’s velocity rather than on the speed of sound. Indeed, this fact enables to concentrate computational power on the waves of interest, see [44], increasing their accuracy.
We refer to Figure 1 to illustrate this phenomenon. Here, we show the numerical solution of a $2\times 2$ linear hyperbolic system with eigenvalues $\lambda_{1}\approx-16$ and $\lambda_{2}\approx 1$. We use a first-order backward Euler method with two different time-steps. First, we fix a time-step $\Delta t_{\text{fast}}$ based on the fast wave, resulting in the dashed red line, and then we choose the time-step $\Delta t_{\text{slow}}$ based on the slow wave, obtaining the solid blue line. It is apparent that the solution with $\Delta t_{\text{fast}}$ (which is the one an explicit scheme would be forced to use) is more diffusive on the slow wave, although it captures the fast wave. On the other hand, the solution computed with $\Delta t_{\text{slow}}$, which can only be used by an implicit scheme, is more accurate on the slow wave. Thus, implicit time stepping allows to choose which waves will be better resolved by the scheme.
@@
Although implicit time-integration methods can have superior stability properties, they require the solution of nonlinear algebraic equations at each time-step, which can be computationally intensive. Therefore, the choice between explicit and implicit schemes depends on the specific problem being solved and the desired trade-off between computational efficiency and accuracy.
@@
In this paper, first of all, we analyze dissipation and dispersion properties of DIRK-DG schemes. Dissipation-dispersion relations are essential tools for understanding how accurately the method captures wave propagation. Wave analysis helps in designing schemes that minimize numerical artifacts such as spurious oscillations and excessive damping. The dissipation-dispersion properties of numerical methods have been extensively studied, including DG methods. However, much of this research has focused on semi-discrete DG schemes [2, 3, 32, 31] or fully-discrete explicit RKDG methods [59, 29].
@@
We mention the following works on the implicit integration of hyperbolic systems. A similar idea of [43, 44] was introduced in [28] where an explicit predictor was used within a fifth-order implicit WENO scheme. Fully nonlinear implicit scheme have been more investigated. For instance, a third order RADAU time integrator combined with a third-order WENO reconstruction can be found in [6]. Fully implicit approaches for DG methods have been considered in [50, 51]. Semi-implicit, implicit-explicit, local time-stepping, semi-Lagrangian and active flux treatments of stiff hyperbolic equations were also investigated, e.g., in [20, 26, 27, 8, 9, 13, 42].
The rest of the paper is organized as follows. In Section 2, we provide a comprehensive review of the DG space discretization methods for hyperbolic systems, with a particular emphasis on implicit time-integration techniques. Section 3 presents a detailed Fourier analysis of fully-discrete implicit DG schemes, where we define measures of dissipation and dispersion, focusing specifically on DIRK-DG schemes of second- and third-order. This section also includes analytical expressions for the eigenvalues and relative energies associated with second-order schemes. In Section 4, we introduce the concept of slope limiters in DG schemes and present the novel framework aimed at reducing nonlinearity by precomputing limiters using first-order implicit predictors. Finally, Section 5 showcases numerical simulations for both scalar equations and systems, demonstrating the practical application and effectiveness of the proposed methods, which reflects the insights of the dissipation-dispersion analysis. We conclude the paper in Section 6 by summarizing the results and discussing future perspectives.
@@ 2.1 Space approximation
$\{\zeta_{0}(x)=1\},$ for $\quad\text{for }\ p=0$ (2.8)
$\{\zeta_{0}(x),\zeta_{1}(x)=x\},$ for $\quad\text{for }\ p=1$ (2.9)
$\{\zeta_{0}(x),\zeta_{1}(x),\zeta_{2}(x)=\frac{1}{2}(3x^{2}-1)\},$ for $\quad\text{for }\ p=2$ (2.10)
@@ 2.2 Implicit time-integration
$\mathbf{u}^{(l),(i)}_{j}=\mathbf{u}^{(l),n}_{j}-(2l+1)\frac{\Delta t}{h}\sum_{\kappa=1}^{i}a_{i\kappa}K_{j}^{(\kappa)}.$ (2.13)
@@
$\mathbf{u}_{j}^{(l),(i)}+(2l+1)\frac{a_{ii}\Delta t}{h}K_{j}^{(i)}-\mathbf{u}_{j}^{(l),n}+(2l+1)\frac{\Delta t}{h}\sum_{\kappa=1}^{i-1}a_{ki}K_{j}^{(\kappa)}=\mathbf{0},\quad j=1,\dots,N,\ l=0,\dots,p.$ (2.14)
Above we have highlighted the term $K^{(i)}$, which makes the system for the $i$-th stage value nonlinear. In fact, the other space approximation terms $K^{(\kappa)}$, $\kappa=1,\dots,i-1$, are already available, thanks to the structure of DIRK schemes. Observe that the stencil of (2.14) is given by the cells $j-1$, $j$ and $j+1$, independently of the order of the DG approximation. Moreover, up to now, the non-linearity of (2.14) is only due to the nonlinear flux function $\mathbf{f}$, which enters in the computation of the numerical fluxes and of the quadrature rule.
@@ 3 Dissipation-dispersion analysis of DIRK-DG schemes
which represents a sinusoidal wave train with a wave number $\mathrm{k}$ and a frequency $\omega=a\mathrm{k}$. In order to compare numerical and exact solutions, we consider the projection $\Pi_{V_{h}^{p}}$ of (3.2) to the DG space $V_{h}^{p}$, so that we can write
@@
where $\omega\in\mathbb{C}$ is a function of the wave number $\mathrm{k}$ and defines the numerical frequency of the scheme. We should expect that the numerical frequency matches $\mathrm{k}a$ to recover the projection of the exact solution. Naively, one could expect that all modes contribute in the same way to approximate $\omega=\mathrm{k}a$ $\forall\,l=0,\dots,p$. We will see, however, that the situation is far more complicated.
@@ 3.1 Numerical dissipation and dispersion of fully-discrete DG schemes
[…]
@@
where $D^{\pm}$, $D$ are $p\times p$ time-independent matrices whose elements are
@@
$\omega_{1}=-i\mathrm{k}-\frac{\mathrm{k}^{4}}{72}h^{3}-\frac{i\mathrm{k}^{5}}{270}h^{4}+O(h^{5})$ and $\omega_{2}=-\frac{6}{h}+O(1)$;
@@
$\omega_{1}=-i\mathrm{k}-\frac{\mathrm{k}^{6}}{7200}h^{5}-\frac{i\mathrm{k}^{7}}{42000}h^{6}+O(h^{7})$ and $\omega_{2,3}=\frac{-3\pm\sqrt{51}i}{h}+O(1)$.
This proves that $A$ is diagonalizable with $p+1$, $k=1,2$, distinct eigenvalues. One of these eigenvalues is physically relevant since it approximates $-i\mathrm{k}$ with dissipation error of order $12p+$ 1 and dispersion error of order $22p+$ 2.
@@
$\sum_{l=0}^{p}w_{l}\boldsymbol{\theta_{l}}=\boldsymbol{\mu},$ (3.12)
@@
$\mathbf{W}^{n}=M^{n}\boldsymbol{\mu}=\sum_{l=0}^{p}w_{l}\,M^{n}\,\boldsymbol{\theta}_{l}=\sum_{l=0}^{p}w_{l}\,\lambda^{n}_{l}\,\boldsymbol{\theta}_{l},$ (3.13)
which is a superposition of eigenmodes along the eigenvectors of $M$ weighted by the expansion coefficients $\{w_{l}\}_{l=0}^{p}$ of the initial condition, and where $\{\lambda_{l}\}_{l=0}^{p}$ are the eigenvalues of $M$. Recalling Ansatz (3.9), we have that the numerical solution of the degrees of freedom $\mathbf{U}_{j}$ at time $t^{n}$ is
$\mathbf{U}_{j}^{n}=\sum_{l=0}^{p}w_{l}\,\lambda^{n}_{l}\,\boldsymbol{\theta}_{l}\,e^{i\mathrm{k}x_{j}}\quad\Rightarrow\quad u^{(m),n}_{j}=\sum_{l=0}^{p}w_{l}\,\lambda^{n}_{l}\theta_{ml}\,e^{i\mathrm{k}x_{j}}\quad m=0,\ldots,p,$ (3.14)
@@
$u_{h}(x,t^{n})=\sum_{m=0}^{p}\left(\sum_{l=0}^{p}w_{l}\,\lambda^{n}_{l}\theta_{ml}\,e^{i\mathrm{k}x_{j}}\right)v^{(m)}_{j}(x)=\left\langle\mathbf{U}_{j}^{n},\mathbf{v}_{j}(x)\right\rangle=\left\langle\mathbf{W}^{n},\mathbf{v}_{j}(x)\right\rangle e^{i\mathrm{k}x_{j}},$ (3.15)
@@
$\Pi_{V_{h}^{p}}(e^{i\mathrm{k}(x-at)})=\sum_{m=0}^{p}\left(\sum_{l=0}^{p}w_{l}\theta_{ml}e^{i\mathrm{k}(x_{j}-at^{n})}\right)v^{(m)}_{j}(x)=\left\langle\sum_{l=0}^{p}w_{l}e^{-i\mathrm{k}at^{n}}\boldsymbol{\theta}_{l},\mathbf{v}_{j}(x)\right\rangle e^{i\mathrm{k}x_{j}}.$ (3.16)
The relations for dissipation and dispersion introduced by the numerical scheme are determined from the absolute error between (3.15) and (3.16):
$\mathcal{E}^{n}=\sum_{l=0}^{p}w_{l}\boldsymbol{\theta}_{l}\left(\lambda_{l}^{n}-e^{-i\mathrm{k}at^{n}}\right),$ (3.17)
which shows that each eigenmode imparts a specific level of dissipation and dispersion to the numerical solution. Notice that in a single time-step the numerical scheme neither dissipates nor disperses if either $\lambda_{l}=e^{-i\mathrm{k}a\Delta t}$  $\forall\,l=0,\dots,p$, or there exists $\hat{l}\in\{0,\dots,p\}$ such that $\lambda_{\hat{l}}=e^{-i\mathrm{k}a\Delta t}$ and $w_{l}=0$  $\forall\,l\neq\hat{l}$.
@@
Thus, we observe that the numerical dissipation in an eigenmode is characterized by $\mathfrak{Im}(\Omega_{l})<0$, whereas numerical dispersion is present if $\mathfrak{Re}(\Omega_{l})-aK\neq 0$. To measure how much a single eigenmode contributes to the numerical dissipation and dispersion, we observe that
$\left\lVert\mathcal{E}^{n}\right\rVert^{2}\leq\sum_{l=0}^{p}|w_{l}|^{2}|e^{-i(\Omega_{l}-aK)n\Delta t}-1|^{2},$ (3.19)
@@
$\beta_{l}=\frac{|w_{l}|^{2}}{\sum_{m=0}^{p}|w_{m}|^{2}},\quad l=0,\ldots,p,$ (3.20)

### v3

@@ Abstract
The application of discontinuous Galerkin (DG) schemes to hyperbolic systems of conservation laws requires a careful interplay between space discretization, carried out with local polynomials and numerical fluxes at inter-cells, and time-integration to yield the final update. An important concern is how the scheme modifies the solution through the notions of numerical dissipation-dispersion. As far as we know, no analysis of these artifacts has been considered for implicit integration of DG methods. The first part of this work intends to fill this gap, showing that the choice of the implicit Runge-Kutta impacts deeply on the quality of the solution. We analyze one-dimensional dissipation-dispersion to select the best combination of the space-time discretization for high Courant numbers.
Then, we apply our findings to the integration of one-dimensional stiff hyperbolic systems. Implicit schemes leverage superior stability properties enabling the selection of time-steps based solely on accuracy requirements. High-order schemes require the introduction of local space limiters which make the whole implicit scheme highly nonlinear. To mitigate the numerical complexity, we propose to use appropriate space limiters that can be precomputed on a first-order prediction of the solution. Numerical experiments explore the performance of this technique on scalar equations and systems.
@@ 1 Introduction
From a numerical viewpoint, stiffness imposes severe restrictions on time-step sizes when using explicit time-integration methods, which significantly increases computational costs. In fact, the Courant-Friedrichs-Lewy (CFL) condition forces the time-step to be small enough to ensure stability, which in turn imposes accuracy on fast-moving waves. To better address the dynamics of the slower-moving fluid, implicit time-integration methods are often used. These methods are unconditionally stable and do not adhere to the CFL condition, allowing to choose the time-step based on the fluid’s velocity rather than on the speed of sound. Indeed, this fact enables to concentrate computational power on the waves of interest, see [44], increasing their accuracy. It is also important to note that, while implicit schemes enable larger time steps and accurate resolution of slow dynamics, this may come at the expense of reduced accuracy for fast acoustic waves when they are present.
We refer to Figure 1 to illustrate, on a simple model problem, the interplay between time-step selection and the accuracy on different wave families. In this toy example, we consider a $2\times 2$ linear hyperbolic system with eigenvalues $\lambda_{1}\approx-16$ and $\lambda_{2}\approx 1$. Using a first-order backward Euler scheme, we compare two choices of the time-step: $\Delta t_{\text{fast}}=h/16$, based on the fast wave (dashed red line), and $\Delta t_{\text{slow}}=h$, based on the slow wave (solid blue line). Here, $h$ represents the mesh-size. As shown, the smaller time-step $\Delta t_{\text{fast}}$ (which an explicit scheme would be constrained to use) provides resolution of the fast wave but is overly diffusive on the slow one. Conversely, the larger step $\Delta t_{\text{slow}}$, feasible only with an implicit scheme, is much more accurate on the slow wave.
@@
This simple illustration highlights a more general principle: implicit time integration enables the user to shift the balance of accuracy towards selected wave families, at the price of solving nonlinear algebraic systems at each step. Hence, while the example is deliberately simplified, the underlying trade-off between stability, accuracy, and computational cost is representative of broader classes of hyperbolic problems.
@@
In this paper, first of all, we analyze one-dimensional dissipation and dispersion properties of DIRK-DG schemes. Dissipation-dispersion relations are essential tools for understanding how accurately the method captures wave propagation. Wave analysis helps in designing schemes that minimize numerical artifacts such as spurious oscillations and excessive damping. The dissipation-dispersion properties of numerical methods have been extensively studied, including DG methods. However, much of this research has focused on semi-discrete DG schemes [2, 3, 32, 31] or fully-discrete explicit RKDG methods [60, 29].
@@
We mention the following works on the implicit integration of hyperbolic systems. A similar idea of [43, 44] was introduced in [28] where an explicit predictor was used within a fifth-order implicit WENO scheme. Fully nonlinear implicit scheme have been more investigated. For instance, a third order RADAU time integrator combined with a third-order WENO reconstruction can be found in [6]. Fully implicit approaches for DG methods have been considered in [51, 52]. We also mention the work [45], where a backward Euler DG scheme combined with a positivity-preserving strategy is shown to be effective in controlling oscillations near shocks. This robustness, however, is mainly due to the strong dissipative nature of backward Euler, which distinguishes it from the low-dissipation high-order DIRK schemes considered in the present study. Semi-implicit, implicit-explicit, local time-stepping, semi-Lagrangian and active flux treatments of stiff hyperbolic equations were also investigated, e.g., in [20, 26, 27, 8, 9, 13, 42].
The rest of the paper is organized as follows. In Section 2, we provide a comprehensive review of the DG space discretization methods for hyperbolic systems, with a particular emphasis on implicit time-integration techniques. Section 3 presents a detailed Fourier analysis of fully-discrete implicit DG schemes, where we define measures of dissipation and dispersion, focusing specifically on DIRK-DG schemes of second- and third-order. This section also includes analytical expressions for the eigenvalues and relative energies associated with second-order schemes. In Section 4, we introduce the concept of slope limiters in DG schemes and present the novel framework aimed at reducing nonlinearity by precomputing limiters using first-order implicit predictors. Finally, Section 5 showcases one-dimensional numerical simulations for both scalar equations and systems, demonstrating the practical application and effectiveness of the proposed methods, which reflects the insights of the dissipation-dispersion analysis. We conclude the paper in Section 6 by summarizing the results and discussing future perspectives.
@@ 2.1 Space approximation
$\{\zeta_{0}(x)=1\},$ $\quad\text{for }\ p=0$ (2.8)
$\{\zeta_{0}(x),\zeta_{1}(x)=x\},$ $\quad\text{for }\ p=1$ (2.9)
$\{\zeta_{0}(x),\zeta_{1}(x),\zeta_{2}(x)=\frac{1}{2}(3x^{2}-1)\},$ $\quad\text{for }\ p=2$ (2.10)
@@ 2.2 Implicit time-integration
$\mathbf{u}^{(l),(i)}_{j}=\mathbf{u}^{(l),n}_{j}-(2l+1)\frac{\Delta t}{h}\sum_{\kappa=1}^{i}a_{i\kappa}K_{j}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(l)},(\kappa)}.$ (2.13)
@@
$\mathbf{u}_{j}^{(l),(i)}+(2l+1)\frac{a_{ii}\Delta t}{h}K_{j}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(l)},(i)}-\mathbf{u}_{j}^{(l),n}+(2l+1)\frac{\Delta t}{h}\sum_{\kappa=1}^{i-1}a_{ki}K_{j}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(l)},(\kappa)}=\mathbf{0},\quad j=1,\dots,N,\ l=0,\dots,p.$ (2.14)
Above we have highlighted the term $K^{(i)}$, which makes the system for the $i$-th stage value nonlinear. In fact, the other space approximation terms $K^{(\kappa)}$, $\kappa=1,\dots,i-1$, are already available, thanks to the structure of DIRK schemes. Observe that the stencil of (2.14) is given by the cells j1j-1, $j$ and $j+1$, independently of the order of the DG approximation. Moreover, up to now, the non-linearity of (2.14) is only due to the nonlinear flux function $\mathbf{f}$, which enters in the computation of the numerical fluxes and of the quadrature rule.
@@ 3 Dissipation-dispersion analysis of DIRK-DG schemes
which represents a sinusoidal wave train with a wave number $\mathrm{k}$ and a frequency $a\mathrm{k}$. In order to compare numerical and exact solutions, we consider the $L^{2}$-projection $\Pi_{V_{h}^{p}}$ of (3.2) to the DG space $V_{h}^{p}$, so that we can write
@@
where $\omega\in\mathbb{C}$ is a function of the wave number $\mathrm{k}$ and defines the numerical frequency of the scheme. We should expect that the numerical frequency matches $\mathrm{k}a$ to recover the projection of the exact solution. Naively, one could expect that all modes contribute in the same way to approximate $\mathrm{k}a$ $\forall\,l=0,\dots,p$. We will see, however, that the situation is far more complicated.
@@ 3.1 Numerical dissipation and dispersion of fully-discrete DG schemes
$\begin{array}[]{c}\displaystyle\frac{h}{2}Z_{l}\,\frac{\mathrm{d}u_{j}^{(l)}(t)}{\mathrm{d}t}-a\sum_{m=0}^{p}Z^{\prime}_{lm}\ u^{(m)}_{j}(t)+a\,\sum_{m=0}^{p}\Big[\left(b^{+}u^{(m)}_{j}(t)+(-1)^{m}b^{-}u^{(m)}_{j+1}(t)\right)\vskip 6.0pt plus 2.0pt minus 2.0pt\\ -(-1)^{l}\left(b^{+}u^{(m)}_{j-1}(t)+(-1)^{m}b^{-}u^{(m)}_{j}(t)\right)\Big]=0,\quad l=0,\ldots,p,\end{array}$ (3.6)
@@
where $^{\pm}, DD$ are $p\times p$ time-independent matrices whose elements are
@@
$e_{1}=-i\mathrm{k}-\frac{\mathrm{k}^{4}}{72}h^{3}-\frac{i\mathrm{k}^{5}}{270}h^{4}+O(h^{5})$ and $e_{2}=-\frac{6}{h}+O(1)$;
@@
$e_{1}=-i\mathrm{k}-\frac{\mathrm{k}^{6}}{7200}h^{5}-\frac{i\mathrm{k}^{7}}{42000}h^{6}+O(h^{7})$ and $e_{2,3}=\frac{-3\pm\sqrt{51}i}{h}+O(1)$.
This proves that $A$ is diagonalizable with $p+1$, $p=1,2$, distinct eigenvalues. One of these eigenvalues is physically relevant since it approximates $i\mathrm{k}$ with dissipation error of order $2p+1$ and dispersion error of order $2p+2$.
@@
$\sum_{l=0}^{p}\nu_{l}\,\boldsymbol{\theta_{l}}=\boldsymbol{\mu},$ (3.12)
@@
$\mathbf{W}^{n}=M^{n}\boldsymbol{\mu}=\sum_{l=0}^{p}\nu_{l}\,M^{n}\,\boldsymbol{\theta}_{l}=\sum_{l=0}^{p}\nu_{l}\,\lambda^{n}_{l}\,\boldsymbol{\theta}_{l},$ (3.13)
which is a superposition of eigenmodes along the eigenvectors of $M$ weighted by the expansion coefficients $\{\nu_{l}\}_{l=0}^{p}$ of the initial condition in (3.12), and where $\{\lambda_{l}\}_{l=0}^{p}$ are the eigenvalues of $M$. Recalling Ansatz (3.9), we have that the numerical solution of the degrees of freedom $\mathbf{U}_{j}$ at time $t^{n}$ is
$\mathbf{U}_{j}^{n}=\sum_{l=0}^{p}\nu_{l}\,\lambda^{n}_{l}\,\boldsymbol{\theta}_{l}\,e^{i\mathrm{k}x_{j}}\quad\Rightarrow\quad u^{(m),n}_{j}=\sum_{l=0}^{p}\nu_{l}\,\lambda^{n}_{l}\theta_{ml}\,e^{i\mathrm{k}x_{j}}\quad m=0,\ldots,p,$ (3.14)
@@
$u_{h}(x,t^{n})=\sum_{m=0}^{p}\left(\sum_{l=0}^{p}\nu_{l}\,\lambda^{n}_{l}\theta_{ml}\,e^{i\mathrm{k}x_{j}}\right)v^{(m)}_{j}(x)=\left\langle\mathbf{U}_{j}^{n},\mathbf{v}_{j}(x)\right\rangle=\left\langle\mathbf{W}^{n},\mathbf{v}_{j}(x)\right\rangle e^{i\mathrm{k}x_{j}},$ (3.15)
where $\mathbf{v}_{j}(x)=(v^{(0)}_{j}(x),\dots,v^{(p)}_{j}(x))^{T}$ and $\langle\cdot,\cdot\rangle$ is the dot-product. Equation (3.15) shows that $\mathbf{U}_{j}^{n}=\mathbf{W}^{n}e^{i\mathrm{k}x_{j}}$ is indeed a solution of the DG system, with $\mathbf{W}^{n}$ not depending on the cell.
$\Pi_{V_{h}^{p}}(e^{i\mathrm{k}(x-at)})=\sum_{m=0}^{p}\left(\sum_{l=0}^{p}\nu_{l}\theta_{ml}e^{i\mathrm{k}(x_{j}-at^{n})}\right)v^{(m)}_{j}(x)=\left\langle\sum_{l=0}^{p}\nu_{l}e^{-i\mathrm{k}at^{n}}\boldsymbol{\theta}_{l},\mathbf{v}_{j}(x)\right\rangle e^{i\mathrm{k}x_{j}}.$ (3.16)
The relations for dissipation and dispersion introduced by the numerical scheme are determined from the absolute error between (3.15) and (3.16).
Proposition 3.2.
Let $M$ be the $(p+1)\times(p+1)$ matrix representing the fully-discrete numerical scheme applied to the linear problem (3.1), with exact solution (3.2). Let $M$ be a diagonalizable matrix, and let $\{(\lambda_{l},\boldsymbol{\theta}_{l})\}_{l=0}^{p}$ denote the set of eigenvalues of $M$ and their corresponding normalized, linearly independent eigenvectors. At a fixed time $t^{n}=n\Delta t$, the error between the approximate solution (2.3) and the exact solution, is defined by the following components,
$\mathcal{E}^{n}_{m}=\sum_{l=0}^{p}\nu_{l}\,\boldsymbol{\theta}_{ml}\left(\lambda_{l}^{n}-e^{-i\mathrm{k}at^{n}}\right),\,m=0,\ldots,p,$ (3.17)
where the coefficients $\{\nu_{l}\}_{l=0}^{p}$ are defined in (3.12) with $\boldsymbol{\mu}(\mathrm{k}h)=\left(c_{0}(\mathrm{k}h),\ldots,c_{p}(\mathrm{k}h)\right)^{T}$, $c_{l}(\mathrm{k}h)$, $l=0,\ldots,p$, given in (3.4).
Proof.
Since $M$ is diagonalizable, the approximate solution (2.3) can be expressed as in (3.15). Then, (3.17) is obtained by subtracting (3.15) and (3.16). ∎
@@
Thus, we observe that the numerical dissipation in an eigenmode is characterized by $\mathfrak{Im}(\Omega_{l})<0$, whereas numerical dispersion is present if $\mathfrak{Re}(\Omega_{l})-aK\neq 0$. To measure how much a single eigenmode contributes to the numerical dissipation and dispersion, we observe that, for each $m=0,\ldots,p$
$\left\lVert\mathcal{E}^{n}_{m}\right\rVert^{2}\leq\sum_{l=0}^{p}|\nu_{l}|^{2}|e^{-i(\Omega_{l}-aK)n\Delta t}-1|^{2},$ (3.19)
@@
$\beta_{l}=\frac{|\nu_{l}|^{2}}{\sum_{m=0}^{p}|\nu_{m}|^{2}},\quad l=0,\ldots,p,$ (3.20)